if transaction is saying , On 1 november 2020 old equipment with an original cost price of R280 000 and a carrying value of R154 000 on 1 march 2020 was traded in against new equipment costing R540 000.after the trade in BONA traders still owed the creditors R410 000.
15 %on cost ..our financial year ends at 28 february 2021...help me understand the question
1
Expert's answer
2021-06-24T12:42:40-0400

Solution:

The question is about the various costs applied to an asset, how it is valued, depreciation and accumulated depreciation determination, and how old assets can be traded in with new assets.

According to the values given, we can determine the following values:

Carrying value = Original cost price – accumulated depreciation

Accumulated depreciation = original cost price – carrying value

Accumulated depreciation = 280,000 – 154,000 = 126,000

Depreciation = 15"\%" of cost = 15% "\times" 280,000 = 42,000 per year


Period of the old asset = "\frac{126,000}{42,000} = 3 \;years". This means that the asset was purchased 3 years ago or has been in operation for the last 3 years.

Balance owed = Cost of new equipment – the fair market value of old equipment

The fair market value of old equipment = Cost of new equipment – balance owed

Fair market value of old equipment = 540,000 – 410,000 = 130,000